# EnrichEvent: Enriching Social Data with Contextual Information for Emerging Event Extraction

### v4

@@ 5.4 Evaluation of Trending Data Extraction
(2) Efficacy: To evaluate the performance of the trend detection model, we measure commonly used classification metrics, e.g., Precision, Recall, and F1-Score. We demonstrate the results of the trend detection model in Table 3 and draw some key findings as follows:
@@
Table 3: Performance of trend detection model
@@
6 Conclusion
Event detection on social media platforms is an active research area, and identification of unspecified events can provide beneficial awareness for making crucial decisions in different fields. In this paper, we design a novel event detection framework to identify unspecified events from social media data streams. Our proposed language-independent framework can easily be adjusted for different data types like news, Wikipedia, etc. In this work, we utilize combinations of lexical and contextual knowledge of tweets. The contextual knowledge enriches the model’s perspective and assists the model in understanding the relationships between tweets more sufficiently. We empirically demonstrate the superiority of our framework in detecting and distinguishing social events compared to the baselines through conducted experiments. An intriguing future research direction would be leveraging LLMs for event summarization and enhancing the process of generating event chains.
Appendix A Appendix
A.1 Datasets Details
Labeling of Trend Detection Dataset: To prepare a generalized trend detection dataset, we defined 12 categories for tweets’ topics. Furthermore, we explored various accounts and collected all the posted tweets from all the experts and active users in each category using TwitterAPI. We leveraged domain-specific key phrases to label the collected tweets in each category. Specifically, we defined a collection of key phrases for each category and checked their occurrence within the tweets to label the dataset. We consider labeling one for the tweets even if one of the key phrases appears in them.
Table 4: 12 categories for tweets’ topics

### v5

@@ V-D Evaluation of Trending Data Extraction
(2) Efficacy: To evaluate the performance of the trend detection model, we measure commonly used classification metrics, e.g., Precision, Recall, and F1-Score. We demonstrate the results of the trend detection model in Table III and draw some key findings as follows:
@@
TABLE III: Performance of trend detection model
@@
VI Conclusion
Event detection on social media platforms is an active research area, and identification of unspecified events can provide beneficial awareness for making crucial decisions in different fields. In this paper, we design a novel event detection framework to identify unspecified events from social media data streams. Our proposed language-independent framework can easily be adjusted for different data types like news, Wikipedia, etc. In this work, we utilize combinations of lexical and contextual knowledge of tweets. Contextual knowledge enriches the model’s perspective and helps the model understand the relationships between tweets more sufficiently. We empirically demonstrate the superiority of our framework in detecting and distinguishing social events compared to the baselines through conducted experiments. An intriguing future research direction would be leveraging LLMs for event summarization and enhancing the process of generating event chains.
[EnrichEvent Output] After detecting the events, the EnrichEvent framework returns a JSON file for each event as an output, including detailed information about the detected events, which is listed in Table IV. This information contains the summary of each event, the period in which each event occurred, and detailed data about the hashtags, users, and named entities that referred to each event.
TABLE IV: A summary of the existing information about the events in EnrichEvent’s Output.
Key Description
Event ID A unique identifier assigned to the event.
Event Period The period during which the event occurs.
Event Summary A brief description of the event.
Tweet with Most Likes The tweet that received the highest number of likes.
Tweet with Most Retweets The tweet that was retweeted the most.
Tweet with Most Replies The tweet with the highest number of replies.
Count Unique Hashtags The total number of distinct hashtags used.
Count Total Hashtags The aggregate number of all hashtags used.
Most Frequent Hashtags The hashtags that appear most frequently in the event.
Count Unique Users The total count of distinct users involved in the event.
Most Frequent Users The users who posted the most tweets.
Count Unique Entities The total count of distinct entities mentioned in the event.
Count Total Entities The aggregate count of all entities mentioned.
Most Frequent Entities The entities mentioned most frequently in the event.
Word Cloud A visual representation of the most common words related to the event.
[Datasets Details] Labeling of Trend Detection Dataset: To prepare a generalized trend detection dataset, we defined 12 categories for tweets’ topics. Furthermore, we explored various accounts and collected all the tweets posted by experts and active users in each category using TwitterAPI. We leveraged domain-specific key phrases to label the collected tweets in each category. Specifically, we defined a collection of key phrases for each category and checked their occurrence within the tweets to label the dataset. We consider labeling one for the tweets even if one of the key phrases appears in them.
TABLE V: Proposed categories for tweets’ topics. We aimed to collect adequate tweets in each category and train our proposed trend detection model in various fields.